LASER VAPORIZED SPECTRA OF A NEW RED SYSTEM OF SiN

K. S. OJHA and \underline{R. GOPAL}, Laser and Spectroscopy Laboratory, Department of Physics, University of Allahabad, ALLAHABAD - 211002, INDIA.

The spectrum has been recorded for first time in the region 575 - 940 nm along with known blue-green systems using second harmonic of pulsed Nd: YAG laser for vaporization of rotating SiN target inside the ablation chamber and computer controlled Spex TRIAX 320 m monochromator fitted with TE cooled ICCD as a detector. It consists of 50 new bands assigned to Δv = 0, ±1, ±2, −3, −4 and −5 sequences a new band system with (0,0) band lying at 787 nm. The lower state vibrational constants are same as that of A^{2}Π state reported by earlier workers. The vibrational constants of upper excited state are as follows:

v_{e} = 12752.4 cm^{−1} ω′_{e} = 945.3 cm^{−1} ω′_{e}x′_{e} = 6.20 cm^{−1}
ω′′_{e} = 1044.4 cm^{−1} ω′′_{e}x′′_{e} = 6.26 cm^{−1}